# ExoJAX Retrievals of VLT/CRIRES Spectra of Luhman 16AB: C/O Ratios and Systematic Uncertainties

### v1

@@ Abstract
facilities: VLT (CRIRES)software: corner (corner), ExoJAX (Kawahara2022; Kawahara2024), FastChem (fastchem; fastchem_cond), JAX (jax2018github), NumPyro (bingham2018pyro; phan2019composable), tinygp (tinygp)
1 Introduction
@@
Detailed characterization of brown dwarfs and exoplanets is becoming possible with high-resolution spectroscopy. In addition to molecular detections with the cross-correlation function, retrieval analyses that directly model spectra to derive molecular abundances and temperature-pressure (TTPP) structures are now commonly applied (e.g., Birkby2018; Snellen2025). Atmospheric properties derived from such retrieval analyses provide valuable insights into the formation pathways of giant planets and brown dwarfs. For example, the relationship between a gas giant’s C/O ratio and its formation environment has long been discussed as an important diagnostic: generally, star-like formation is expected to inherit the elemental ratios of the natal cloud, whereas core accretion within a disk — where solids and gas evolve separately — can alter the initial composition depending on the formation location (e.g., Oberg2011; Espinoza2023). Empirically, C/O ratios measured for wide-orbit gas giants/super-Jupiters from emission spectroscopy have tended to be roughly solar, whereas those inferred for hot Jupiters from transmission spectroscopy, including with JWST, exhibit a wider dispersion, potentially reflecting differences in formation processes (e.g., Walsh2025). While this study focuses on the C/O ratio, elemental abundances such as [C/H] can also provide complementary diagnostics for distinguishing between planetary and brown dwarf formation pathways (e.g., Xuan2024a; Wang2025).
@@
To advance these discussions, brown dwarfs are important comparison targets, providing anchor points in addition to gas giants: while the formation of cold Jupiters remains debated, brown dwarfs likely form in a star-like manner and thus can provide a baseline for C/O. However, the number of available high-resolution measurements remains limited (Figure 1). Since brown dwarfs are often observable at higher S/N than planetary-mass objects, they also offer valuable testbeds for assessing retrieval systematics. Despite rapid progress of atmospheric retrieval, high-resolution retrievals still carry systematic uncertainties from cloud modeling, molecular line lists, TTPP parameterizations, and chemical–equilibrium assumptions, and a community standard has yet to emerge. For example, Picos2024 compared spectral models built on two distinct gradient assumptions of TTPP profiles and showed that the inferred surface gravity can shift appreciably with the chosen TTPP model, while deRegt2024 demonstrated that using alternative $\mathrm{H_{2}O}$ line lists leads to different inferences for the isotopologue ratio $^{12}\mathrm{CO}/^{13}\mathrm{CO}$. Testing their impact with high-quality data is important both for comparative planetology and for understanding these processes themselves.
@@
The paper is organized as follows. Section 2 describes the data, and Section 3 outlines the modeling framework used for the retrievals. In Section 4, we analyze the line-list dependence within a power-law TTPP framework. Section 5 then introduces a more flexible Gaussian-process TTPP profile, demonstrates its utility, and discusses how assumptions about the TTPP structure propagate into other parameters and their interpretation. Section 6 summarizes our findings, discusses implications, and outlines directions for future work.
@@ 2 Data
3 Retrieval Framework
3.1 Spectrum Model
@@
We modeled the high-resolution spectra of Luhman 16AB with the ExoJAX spectral framework (Kawahara2022; Kawahara2024). ExoJAX solves the radiative transfer by computing, for each pressure interval $\Delta P$ between adjacent levels, the optical–depth increment as the sum of the molecular contribution in a given layer $\sum_{i}\Delta\tau_{\mathrm{mol},i}$, collision–induced absorption (CIA) $\Delta\tau_{\mathrm{CIA}}$, and cloud opacity $\Delta\tau_{\mathrm{c}}$:
@@
where $C$ is the (dimensionless) column optical depth of the cloud. This functional form is motivated by the vertical cloud opacity profiles in the Sonora diamondback models (Morley2024). Given the narrow wavelength range analyzed here, we neglect any wavelength dependence of the cloud optical depth. Because $w_{\mathrm{c}}$ and $P_{\mathrm{c}}$ are strongly degenerate, we fix $w_{\mathrm{c}}=0.3$. To emulate an effectively optically thick gray deck, we fix the column optical depth to a large arbitrary constant, $C=500$.
Using the resulting $\Delta\tau(\nu)$, we solve the radiative transfer with an eight-stream approximation to obtain the local one-dimensional emergent spectrum $S_{\mathrm{raw}}(\nu)$. We then convolve with a rotational broadening kernel computed from the limb-darkening law and the projected rotational velocity $V\sin i$ to obtain the disk-integrated spectrum $F_{\mathrm{raw}}(\nu)$. For limb darkening, we adopt the quadratic law with the $(q_{1},q_{2})$ parametrization as proposed by Kipping2013.
@@
3.2 Inference
ExoJAX is implemented with JAX (jax2018github), a high-performance Python library that supports GPU acceleration and automatic differentiation. Consequently, our spectral model can be evaluated on GPUs and differentiated with respect to arbitrary parameters, enabling gradient-based inference.
Leveraging these features, we infer posterior distributions of the model parameters with Hamiltonian Monte Carlo (HMC; HMC), specifically the No-U-Turn Sampler (NUTS; NUTS) implemented in NumPyro. With prior $p(\boldsymbol{\theta})$, the posterior is given by Bayes’ theorem:
@@ 3.2 Inference
To account for correlated noise, we use a Gaussian-process (GP) likelihood (Kawahara2022),
@@
The set of inferred parameters and their priors is listed in Table 1. For $\log g$, we imposed Normal priors informed by external constraints. Using the system age of $510\pm 95\,\mathrm{Myr}$ from gagne2023 together with the dynamical masses $M_{A}=33.5\pm 0.3\,\mathrm{M_{J}}$ and $M_{B}=28.6\pm 0.3\,\mathrm{M_{J}}$ from lazorenko2018, we interpolated the brown dwarf evolutionary model of Baraffe2003 to obtain surface gravities. Since the surface gravity is given by $g=GM/R^{2}$ with brown dwarf mass $M$, radius $R$, and the gravitational constant $G$, we propagated the quoted mass uncertainties ($\sim 1\%$) and assumed a 10% fractional uncertainty on the radius to set the prior widths, yielding Normal priors $\log g_{A}=4.93\pm 0.09$ and $\log g_{B}=4.86\pm 0.09$, each truncated at $3\sigma$. Sampling used HMC–NUTS with a dense mass matrix. For numerical stability and mixing, parameters were affinely re-scaled so that the prior support maps to $\mathcal{O}(1)$ in the sampler’s internal coordinates. Sampling/adaptation were performed in the transformed space and results are reported after back-transformation.
4 Retrievals Using a Power-Law TTPP Profile
@@
4.1 Line Lists
High-resolution retrievals for brown dwarfs commonly use line lists from ExoMol and HITEMP. ExoMol provides theoretically computed line data tailored to brown dwarf and exoplanet atmospheres, assuming an $\mathrm{H_{2}}$$\mathrm{He}$ background. HITEMP, by contrast, is based largely on high-temperature laboratory measurements and typically adopts an Earth background (air) for pressure broadening. More recently, an $\mathrm{H_{2}}$$\mathrm{He}$ broadened variant of HITEMP has become available, obtained by converting the air-broadening parameters; the conversion employs semi-empirical models anchored to validated laboratory measurements and theoretical predictions (tan2022).
4.1.1 ExoMol
According to ExoMol2016, ExoMol supplies transition wavenumbers $\hat{\nu}$, Einstein coefficients $A$, state energies $E$, and statistical weights $\mathsf{g}$, from which the line strength $S(T)$ can be computed as
@@ 4.1.1 ExoMol
4.1.2 HITEMP
According to HiTEMP, HITEMP provides $\hat{\nu}$, $S(T_{\mathrm{ref}})$ at $T_{\mathrm{ref}}=296\,\mathrm{K}$, and $E_{\mathrm{low}}$, from which
@@ 4.1.2 HITEMP
with $P_{\mathrm{ref}}=1\,\mathrm{atm}$. For some molecules, $\mathrm{H_{2}}$-broadened parameters $\gamma_{\mathrm{ref,,H_{2}}}$ and $\mathsf{n}_{\mathrm{H_{2}}}$ are available, allowing $\mathrm{H_{2}}$-background pressure broadening to be computed via Eq. (18) (tan2022). In what follows, we refer to air-broadened HITEMP as $\mathrm{HITEMP_{Air}}$ and $\mathrm{H_{2}}$-broadened HITEMP as $\mathrm{HITEMP_{H_{2}}}$. Nominally, air broadening parameters are not appropriate for brown dwarf atmospheres. However, the laboratory experiment by 2025ApJ...984...92H suggests that these parameters are not necessarily incorrect, so we include them in the comparison.
4.1.3 Adopted Line Lists in This Study
Given that brown dwarf atmospheres are dominated by $\mathrm{H_{2}}$, we adopt the ExoMol POKAZATEL for $\mathrm{H_{2}O}$ (POKAZATEL), ExoMol MM for $\mathrm{CH_{4}}$ (MM), and the ExoMol Coxon–Hajig line list for $\mathrm{HF}$ (Coxon-Hajig). To assess the impact of the CO line list — which exhibits prominent absorption in the $K$ band — on the retrieval results, we consider the following three CO line list configurations:
@@ 4.1.3 Adopted Line Lists in This Study
ExoMol CO line list Li2015 (Li2015), assuming an $\mathrm{H_{2}}$ background.
@@
HITEMP CO line list (HITEMPCO) with air-broadening parameters.
@@
An $\mathrm{H_{2}}$-broadened version of HITEMP CO based on the semi-empirical models constructed on available and validated experimental measurements and theoretical predictions (tan2022).
For all three CO line lists, the same Voigt profile treatment is applied, so that the differences in the retrieval results originate only from the broadening parameters. The selection of molecular species was referenced from the $K$-band high-resolution spectral analysis of Luhman 16AB with IGRINS/Gemini South reported by Ishikawa2025. We also searched for the isotopologue $^{13}\mathrm{CO}$, but it was not detected in the present analysis. This is consistent with Ishikawa2025, who did not detect $^{13}\mathrm{CO}$ in the same wavelength region. Therefore, we did not include $^{13}\mathrm{CO}$ opacity in the spectral model.
4.2 Results
@@
Figure 4: Retrieved TTPP profiles and normalized contribution fuctions for Luhman 16A (left) and Luhman 16B (right) under the $\mathrm{HITEMP_{H_{2}}\,\,CO}$ configuration. For each object, the left panel shows the contribution function, and the right panel shows the posterior TTPP profile. The horizontal shaded band marks the inferred cloud deck pressure, $\log_{10}P\left(\tau_{c}=1\right)$. Note that $\tau_{c}$ denotes the cumulative optical depth due only to cloud opacity.
Blue, purple, and red lines are Sonora Diamondback (SD) TTPP models (Morley2024). Black curves in the right panels indicate condensation curves for the principal brown dwarf cloud species identified by Morley2024 (condensation curve formulas from Visscher2010; Wakeford2017).
@@
The retrieved TTPP profiles under the power-law + gray-cloud + $\mathrm{HITEMP_{H_{2}}\,\,CO}$ configuration are shown for each object in Figure 4. Across all line lists, the inferred temperature ranges in the regions of largest contribution are consistent with the Sonora diamondback models (Morley2024) at the literature effective temperatures. Here, $\tau_{c}$ denotes the cumulative optical depth due only to the gray-cloud opacity integrated downward from the top of the atmosphere, and the pressure level at $\tau_{c}=1$ corresponds to the cloud deck in our model. The inferred cloud-deck pressures fall within the altitude ranges predicted by the condensation curves in the Sonora Diamondback models, indicating broad physical consistency with the cloud-base pressures of plausible condensates, although we do not attempt to identify a specific cloud species.
@@
where the solar ratios $\left(\mathrm{X/H}\right)_{\odot}$ follow Asplund2021. For both Luhman 16A and 16B, the VMR-inferred C/O is $\sim 0.67$ — slightly above solar (Figure 3). We nevertheless find that the inferred C/O shifts by $\sim 7\%$ depending on the adopted line list. Differences in the line shapes alter the inferred cloud-deck pressure and, in turn, the relative VMRs.
To interpret the retrieved abundances, we computed thermochemical-equilibrium VMRs with FastChem including equilibrium condensation (fastchem; fastchem_cond). In these calculations, we adjusted the solar C, O, and F abundances of Asplund2021 by our retrieved $\mathrm{[X/H]}$ values (all other elements fixed to solar). Asplund2021 tabulate $A(\mathrm{X})\equiv\log_{10}(n_{\mathrm{X}}/n_{\mathrm{H}})+12$ with $A(\mathrm{H})=12$; accordingly we adopt
@@
evaluated separately for Luhman 16A and 16B using their respective $\mathrm{[X/H]}$. Numerically, $A_{\odot}(\mathrm{C})=8.46$, $A_{\odot}(\mathrm{O})=8.69$, and $A_{\odot}(\mathrm{F})=4.40$ (Asplund2021). The equilibrium VMRs computed with these inputs and the retrieved TTPP profiles (Figure 5) agree well for $\mathrm{CO}$, $\mathrm{H_{2}O}$, and $\mathrm{HF}$, indicating compatibility with chemical equilibrium under the inferred thermal structures. By contrast, the retrieved $\mathrm{CH_{4}}$ VMRs for both objects are about an order of magnitude lower than the FastChem predictions; we revisit this discrepancy in Section 5.3.
@@
4.2.1 Comparison with Medium-Resolution Flux Density

### v2

@@ Abstract
facilities: VLT (CRIRES)software: corner (Foreman-Mackey, 2016), ExoJAX (Kawahara et al., 2022, 2025), FastChem (Stock et al., 2018; Kitzmann et al., 2024), JAX (Bradbury et al., 2018), NumPyro (Bingham et al., 2018; Phan et al., 2019), tinygp (Foreman-Mackey et al., 2024)
I Introduction
@@
Detailed characterization of brown dwarfs and exoplanets is becoming possible with high-resolution spectroscopy. In addition to molecular detections with the cross-correlation function, retrieval analyses that directly model spectra to derive molecular abundances and temperature-pressure (TTPP) structures are now commonly applied (e.g., Birkby, 2018; Snellen, 2025). Atmospheric properties derived from such retrieval analyses provide valuable insights into the formation pathways of giant planets and brown dwarfs. For example, the relationship between a gas giant’s C/O ratio and its formation environment has long been discussed as an important diagnostic: generally, star-like formation is expected to inherit the elemental ratios of the natal cloud, whereas core accretion within a disk — where solids and gas evolve separately — can alter the initial composition depending on the formation location (e.g., Öberg et al., 2011). Empirically, C/O ratios measured for wide-orbit gas giants/super-Jupiters from emission spectroscopy have tended to be roughly solar, whereas those inferred for hot Jupiters from transmission spectroscopy, including with JWST, exhibit a wider dispersion, potentially reflecting differences in formation processes (e.g., Walsh, 2025). While this study focuses on the C/O ratio, elemental abundances such as [C/H] can also provide complementary diagnostics for distinguishing between planetary and brown dwarf formation pathways (e.g., Xuan et al., 2024a; Wang, 2025).
@@
To advance these discussions, brown dwarfs are important comparison targets, providing anchor points in addition to gas giants: while the formation of cold Jupiters remains debated, brown dwarfs likely form in a star-like manner and thus can provide a baseline for C/O. However, the number of available high-resolution measurements remains limited (Figure 1). Since brown dwarfs are often observable at higher S/N than planetary-mass objects, they also offer valuable testbeds for assessing retrieval systematics. Despite rapid progress of atmospheric retrieval, high-resolution retrievals still carry systematic uncertainties from cloud modeling, molecular line lists, TTPP parameterizations, and chemical–equilibrium assumptions, and a community standard has yet to emerge. For example, González Picos et al. (2024) compared spectral models built on two distinct gradient assumptions of TTPP profiles and showed that the inferred surface gravity can shift appreciably with the chosen TTPP model, while de Regt et al. (2024) demonstrated that using alternative $\mathrm{H_{2}O}$ line lists leads to different inferences for the isotopologue ratio $^{12}\mathrm{CO}/^{13}\mathrm{CO}$. Testing their impact with high-quality data is important both for comparative planetology and for understanding these processes themselves.
@@
The paper is organized as follows. Section II describes the data, and Section III outlines the modeling framework used for the retrievals. In Section IV, we analyze the line-list dependence within a power-law TTPP framework. Section V then introduces a more flexible Gaussian-process TTPP profile, demonstrates its utility, and discusses how assumptions about the TTPP structure propagate into other parameters and their interpretation. Section VI summarizes our findings, discusses implications, and outlines directions for future work.
@@ II Data
III Retrieval Framework
III.1 Spectrum Model
@@
We modeled the high-resolution spectra of Luhman 16AB with the ExoJAX spectral framework (Kawahara et al., 2022, 2025). ExoJAX solves the radiative transfer by computing, for each pressure interval $\Delta P$ between adjacent levels, the optical–depth increment as the sum of the molecular contribution in a given layer $\sum_{i}\Delta\tau_{\mathrm{mol},i}$, collision–induced absorption (CIA) $\Delta\tau_{\mathrm{CIA}}$, and cloud opacity $\Delta\tau_{\mathrm{c}}$:
@@
where $C$ is the (dimensionless) column optical depth of the cloud. This functional form is motivated by the vertical cloud opacity profiles in the Sonora diamondback models (Morley et al., 2024). Given the narrow wavelength range analyzed here, we neglect any wavelength dependence of the cloud optical depth. Because $w_{\mathrm{c}}$ and $P_{\mathrm{c}}$ are strongly degenerate, we fix $w_{\mathrm{c}}=0.3$. To emulate an effectively optically thick gray deck, we fix the column optical depth to a large arbitrary constant, $C=500$.
Using the resulting $\Delta\tau(\nu)$, we solve the radiative transfer with an eight-stream approximation to obtain the local one-dimensional emergent spectrum $S_{\mathrm{raw}}(\nu)$. We then convolve with a rotational broadening kernel computed from the limb-darkening law and the projected rotational velocity $V\sin i$ to obtain the disk-integrated spectrum $F_{\mathrm{raw}}(\nu)$. For limb darkening, we adopt the quadratic law with the $(q_{1},q_{2})$ parametrization as proposed by Kipping (2013).
@@
III.2 Inference
ExoJAX is implemented with JAX (Bradbury et al., 2018), a high-performance Python library that supports GPU acceleration and automatic differentiation. Consequently, our spectral model can be evaluated on GPUs and differentiated with respect to arbitrary parameters, enabling gradient-based inference.
Leveraging these features, we infer posterior distributions of the model parameters with Hamiltonian Monte Carlo (HMC; Neal, 2011), specifically the No-U-Turn Sampler (NUTS; Hoffman and Gelman, 2014) implemented in NumPyro. With prior $p(\boldsymbol{\theta})$, the posterior is given by Bayes’ theorem:
@@ III.2 Inference
To account for correlated noise, we use a Gaussian-process (GP) likelihood (Kawahara et al., 2022),
@@
The set of inferred parameters and their priors is listed in Table 1. For $\log g$, we imposed Normal priors informed by external constraints. Using the system age of $510\pm 95\,\mathrm{Myr}$ from Gagné et al. (2023) together with the dynamical masses $M_{A}=33.5\pm 0.3\,\mathrm{M_{J}}$ and $M_{B}=28.6\pm 0.3\,\mathrm{M_{J}}$ from Lazorenko and Sahlmann (2018), we interpolated the brown dwarf evolutionary model of Baraffe et al. (2003) to obtain surface gravities. Since the surface gravity is given by $g=GM/R^{2}$ with brown dwarf mass $M$, radius $R$, and the gravitational constant $G$, we propagated the quoted mass uncertainties ($\sim 1\%$) and assumed a 10% fractional uncertainty on the radius to set the prior widths, yielding Normal priors $\log g_{A}=4.93\pm 0.09$ and $\log g_{B}=4.86\pm 0.09$, each truncated at $3\sigma$. Sampling used HMC–NUTS with a dense mass matrix. For numerical stability and mixing, parameters were affinely re-scaled so that the prior support maps to $\mathcal{O}(1)$ in the sampler’s internal coordinates. Sampling/adaptation were performed in the transformed space and results are reported after back-transformation.
IV Retrievals Using a Power-Law TTPP Profile
@@
IV.1 Line Lists
High-resolution retrievals for brown dwarfs commonly use line lists from ExoMol and HITEMP. ExoMol provides theoretically computed line data tailored to brown dwarf and exoplanet atmospheres, assuming an $\mathrm{H_{2}}$$\mathrm{He}$ background. HITEMP, by contrast, is based largely on high-temperature laboratory measurements and typically adopts an Earth background (air) for pressure broadening. More recently, an $\mathrm{H_{2}}$$\mathrm{He}$ broadened variant of HITEMP has become available, obtained by converting the air-broadening parameters; the conversion employs semi-empirical models anchored to validated laboratory measurements and theoretical predictions (Tan et al., 2022).
IV.1.1 ExoMol
According to Tennyson et al. (2016), ExoMol supplies transition wavenumbers $\hat{\nu}$, Einstein coefficients $A$, state energies $E$, and statistical weights $\mathsf{g}$, from which the line strength $S(T)$ can be computed as
@@ IV.1.1 ExoMol
IV.1.2 HITEMP
According to Rothman et al. (2010), HITEMP provides $\hat{\nu}$, $S(T_{\mathrm{ref}})$ at $T_{\mathrm{ref}}=296\,\mathrm{K}$, and $E_{\mathrm{low}}$, from which
@@ IV.1.2 HITEMP
with $P_{\mathrm{ref}}=1\,\mathrm{atm}$. For some molecules, $\mathrm{H_{2}}$-broadened parameters $\gamma_{\mathrm{ref,,H_{2}}}$ and $\mathsf{n}_{\mathrm{H_{2}}}$ are available, allowing $\mathrm{H_{2}}$-background pressure broadening to be computed via Eq. (18) (Tan et al., 2022). In what follows, we refer to air-broadened HITEMP as $\mathrm{HITEMP_{Air}}$ and $\mathrm{H_{2}}$-broadened HITEMP as $\mathrm{HITEMP_{H_{2}}}$. Nominally, air broadening parameters are not appropriate for brown dwarf atmospheres. However, the laboratory experiment by Hosokawa et al. (2025) suggests that these parameters are not necessarily incorrect, so we include them in the comparison.
IV.1.3 Adopted Line Lists in This Study
Given that brown dwarf atmospheres are dominated by $\mathrm{H_{2}}$, we adopt the ExoMol POKAZATEL for $\mathrm{H_{2}O}$ (Polyansky et al., 2018), ExoMol MM for $\mathrm{CH_{4}}$ (Yurchenko et al., 2024), and the ExoMol Coxon–Hajig line list for $\mathrm{HF}$ (Coxon and Hajigeorgiou, 2015). To assess the impact of the CO line list — which exhibits prominent absorption in the $K$ band — on the retrieval results, we consider the following three CO line list configurations:
@@ IV.1.3 Adopted Line Lists in This Study
ExoMol CO line list Li2015 (Li et al., 2015), assuming an $\mathrm{H_{2}}$ background.
@@
HITEMP CO line list (Li et al., 2015) with air-broadening parameters.
@@
An $\mathrm{H_{2}}$-broadened version of HITEMP CO based on the semi-empirical models constructed on available and validated experimental measurements and theoretical predictions (Tan et al., 2022).
For all three CO line lists, the same Voigt profile treatment is applied, so that the differences in the retrieval results originate only from the broadening parameters. The selection of molecular species was referenced from the $K$-band high-resolution spectral analysis of Luhman 16AB with IGRINS/Gemini South reported by Ishikawa et al. (2025). We also searched for the isotopologue $^{13}\mathrm{CO}$, but it was not detected in the present analysis. This is consistent with Ishikawa et al. (2025), who did not detect $^{13}\mathrm{CO}$ in the same wavelength region. Therefore, we did not include $^{13}\mathrm{CO}$ opacity in the spectral model.
IV.2 Results
@@
Figure 4: Retrieved TTPP profiles and normalized contribution fuctions for Luhman 16A (left) and Luhman 16B (right) under the $\mathrm{HITEMP_{H_{2}}\,\,CO}$ configuration. For each object, the left panel shows the contribution function, and the right panel shows the posterior TTPP profile. The horizontal shaded band marks the inferred cloud deck pressure, $\log_{10}P\left(\tau_{c}=1\right)$. Note that $\tau_{c}$ denotes the cumulative optical depth due only to cloud opacity. Blue, purple, and red lines are Sonora Diamondback (SD) TTPP models (Morley et al., 2024). Black curves in the right panels indicate condensation curves for the principal brown dwarf cloud species identified by Morley et al. (2024) (condensation curve formulas from Visscher et al. 2010; Wakeford et al. 2017).
@@
The retrieved TTPP profiles under the power-law + gray-cloud + $\mathrm{HITEMP_{H_{2}}\,\,CO}$ configuration are shown for each object in Figure 4. Across all line lists, the inferred temperature ranges in the regions of largest contribution are consistent with the Sonora diamondback models (Morley et al., 2024) at the literature effective temperatures. Here, $\tau_{c}$ denotes the cumulative optical depth due only to the gray-cloud opacity integrated downward from the top of the atmosphere, and the pressure level at $\tau_{c}=1$ corresponds to the cloud deck in our model. The inferred cloud-deck pressures fall within the altitude ranges predicted by the condensation curves in the Sonora Diamondback models, indicating broad physical consistency with the cloud-base pressures of plausible condensates, although we do not attempt to identify a specific cloud species.
@@
where the solar ratios $\left(\mathrm{X/H}\right)_{\odot}$ follow Asplund et al. (2021). For both Luhman 16A and 16B, the VMR-inferred C/O is $\sim 0.67$ — slightly above solar (Figure 3). We nevertheless find that the inferred C/O shifts by $\sim 7\%$ depending on the adopted line list. Differences in the line shapes alter the inferred cloud-deck pressure and, in turn, the relative VMRs.
To interpret the retrieved abundances, we computed thermochemical-equilibrium VMRs with FastChem including equilibrium condensation (Stock et al., 2018; Kitzmann et al., 2024). In these calculations, we adjusted the solar C, O, and F abundances of Asplund et al. (2021) by our retrieved $\mathrm{[X/H]}$ values (all other elements fixed to solar). Asplund et al. (2021) tabulate $A(\mathrm{X})\equiv\log_{10}(n_{\mathrm{X}}/n_{\mathrm{H}})+12$ with $A(\mathrm{H})=12$; accordingly we adopt
@@
evaluated separately for Luhman 16A and 16B using their respective $\mathrm{[X/H]}$. Numerically, $A_{\odot}(\mathrm{C})=8.46$, $A_{\odot}(\mathrm{O})=8.69$, and $A_{\odot}(\mathrm{F})=4.40$ (Asplund et al., 2021). The equilibrium VMRs computed with these inputs and the retrieved TTPP profiles (Figure 5) agree well for $\mathrm{CO}$, $\mathrm{H_{2}O}$, and $\mathrm{HF}$, indicating compatibility with chemical equilibrium under the inferred thermal structures. By contrast, the retrieved $\mathrm{CH_{4}}$ VMRs for both objects are about an order of magnitude lower than the FastChem predictions; we revisit this discrepancy in Section V.3.
@@
IV.2.1 Comparison with Medium-Resolution Flux Density